Find the measures of two supplementary angles if one angle is 18 degrees more than twice the other.
Answer by nerdybill(7384)

Find the measures of two supplementary angles if one angle is 18 degrees more than twice the other.
.
Let x = measure of one of the angles
then from "18 degrees more than twice the other" we have
2x+18 = measure of second angle
.
Then, because they are "supplementary":
x + 2x+18 = 180
3x+18 = 180
3x = 162
x = 54 degrees
.
Second angle:
2x+18 = 2(54)+18 = 108+18 = 126 degrees
.
answer: 54 and 126 degrees